A bullet is fired at an angle of 30° above the horizontal with a velocity of 600m/s. Find the range the time it returns to the same level and the position and velocity of the bullet 40 s after after it is fired


1
Expert's answer
2021-08-25T17:24:11-0400

Given:

"v_0=600\:\rm m/s"

"\theta=30^{\circ}"

"t_1=40\:\rm s"


The range

"R=\frac{v_0^2\sin2\theta}{g}=\frac{600^2\sin 60^{\circ}}{9.8}=31813\:\rm m"

The time of motion

"t=\frac{2v_0\sin\theta}{g}=\frac{2*600\sin 30^{\circ}}{9.8}=61\:\rm s"

After 40 s

"x=v_0\cos\theta *t_1=600\cos 30^{\circ}*40=20784\:\rm m"

"y=v_0\sin\theta *t_1-\frac{gt_1^2}{2}"

"y=600\sin 30^{\circ}*40-\frac{9.8*40^2}{2}=4160\:\rm m"

The speed of a bullet

"v=\sqrt{(v_0\cos\theta)^2+(v_0\sin\theta-gt_1)}"

"v=\sqrt{(600\cos 30^{\circ})^2+(600\sin30^{\circ}-9.8*40)^2}\\=528\:\rm m/s"